OpenStudy (anonymous):

Jane has $5000. She invests a portion of her money at a simple interest rate of 7% and the rest of her money at a simple interest rate of 6.2%. After one year, the total interest earned on these investments is $318.00. How much money did she invest at %7?

OpenStudy (anonymous):

You need two equations. Let x= the amout earning 7%; Let y= the amount earning 6.2% Since the total invested is $5,000 then x+y=5,000 and the .07x+.062y=318 Rearrange the first equation so that y=5,000-x Substitute into the second equation So you have .07x+(.062)(5,000-x)=318 which simplifies to .008x=8 So x=1,000 dollars